# Bayesian sample size calculations for external validation studies of risk prediction models

### v1

@@ Abstract
Background: Contemporary sample size calculations for external validation of risk prediction models require users to specify fixed (i.e., true) values of assumed model performance metrics alongside target precision levels (e.g., 95% CI widths). However, due to the finite samples of previous studies, our knowledge of true model performance in the target population is uncertain, and so choosing fixed values represents an incomplete picture. As well, for net benefit (NB) as a measure of clinical utility, the relevance of conventional precision-based inference is doubtful. Bayesian approach can address this, by accounting for uncertainty of true model performance and facilitating decision-theoretic metrics including the expected value of sample information (EVSI).
Methods: We propose a general Bayesian algorithm for constructing the joint distribution of predicted risks and response values based on summary statistics of model performance in previous studies. For statistical metrics of performance, we propose sample size determination rules that either target desired expected precision, or a desired assurance probability that the precision criteria will be satisfied. For NB, we propose rules based on optimality assurance (the probability that the planned study correctly identifies the most beneficial strategy) and the EVSI, the expected gain in net benefit from the planned validation study. We showcase these developments in a case study on the validation of a risk prediction model for deterioration of hospitalized COVID-19 patients.
Results: The contemporary approach, based on fixed values of assumed c-statistic, O/E ratio, and calibration slope from in the validation study, with target 95%CI width of, respectively, 0.10, 0.22, and 0.30, would recommend a sample size of 1,056, dictated by the desired precision around the calibration slope. To implement the Bayesian approach and to account for the uncertainty of the model’s predictive performance, we constructed predictive distributions for prevalence, c-statistic, mean calibration error, and calibration slope for a new validation study. Targeting the same expected CI width would result in a sample size of 1,025. However, demanding 90% assurance for meeting these precision criteria would result in a sample size of 1,173. In terms of NB, an optimality assurance of 90% was achieved at a sample size of 306. The EVSI curve demonstrated a diminishing margin beyond sample sizes >500 (<10% gain in expected NB with doubling the sample size).
Conclusion: Compared to the conventional sample size calculation methods, a Bayesian approach requires explicit quantification of uncertainty around model performance, but thereby enables various sample size rules based on expected precision, assurance probabilities, and value of information. In our case study, EVSI calculations indicated that the precision criterion around calibration slope could potentially be relaxed without much utility loss.
1 Background
@@
Developing risk prediction models or validating existing ones in a new population represents significant investment in time, resources, and expertise. Like other empirical experiments, the design of such studies should be based on objective, transparent, and defendable principles. A particular aspect of study design is the sample size of such studies. The field has witnessed significant recent developments on this front1–8. An example is the multi-criteria approach by Riley et al for the sample size required for external validation studies, targeting, targeting pre-specified width of the 95% confidence interval around metrics of model performance2,3. For binary outcomes, such metrics are related to discrimination, calibration, and net benefit (NB). The sample size required for each component is computed separately, with the largest one advertised as the final requirement3.
@@
In addition to target precision criteria, each component of this approach requires as input an assumed true value of the the metric of interest in the target population. In reality, we do not know the true value of model performance metrics with certainty (otherwise there would be no need to conduct the study in the first place). Further, due to sampling variability, the precision obtained in one particular dataset may differ substantially from the expected precision for that sample size. Thus, there is more uncertainty to be accounted for than the Riley criteria allow.
@@
A Bayesian approach towards sample size determination allows uncertainty to be accounted for, which is advantageous on multiple fronts. First, it enables the full use of existing information on model performance at the time the study is designed, rather than forcing the investigator to express their knowledge as the fixed truth. Second, it enables different classes of sample size rules, namely those that target the expected values of precision targets, as well as ‘assurance’-type rules for the probability of meeting (or exceeding) precision targets. This can be insightful as focusing on expected values alone does not guarantee that the desired precision will be achieved in one particular dataset, and the investigator may want a stronger assurance against this. For example, an investigator that desires an interval width of 0.1 for c-statistic might not perceive much benefit from the future interval being narrower, but might have a strong preference against not meeting this target. As such, targeting a sample size that will result in 90% probability that the CI width will be $\leq$ 0.1 might be more appealing than a sample size that targets an expected CI width of 0.1. Finally, when assessing clinical utility, the relevance of precision-based criteria is challenged9,10. A Bayesian approach enables the use of novel, decision-theoretic approaches based on value of information (VoI) analysis, which focus on the expected gain in clinical utility from increasing sample size11.
@@
Bayesian approaches for power and sample size calculations are mostly investigated in the context of experimental studies that are aimed at interrogating a null hypothesis12–15. However, risk model validation studies are not generally hypothesis-driven. Rather, the focus is on the precise estimation of a variety of metrics of model performance. As such, Bayesian sample size considerations in this context is worth exploring. Hence, in this article, we propose a Bayesian version of the sample size formula by Riley et al, focusing on the same metrics of model performance, but enabling the investigator to 1) incorporate their uncertainty around assumed model performance in calculations, 2) use sample size rules that target assurance probabilities for meeting chosen criteria, and 3) use rules based on VoI analysis for clinical utility. The proposed framework can be used in two general ways: to quantify the anticipated precision or VoI outcomes from a planned study if the sample size is fixed (as is the case with validation studies that are based on already collected data), or to determine the minimum sample size that achieves pre-specified precision or VoI criteria. This framework can also be used in a hybrid form: determine the sample size based on certain criteria (e.g., targeting CI width for c-statistic and calibration slope), and investigate the consequence of the chosen sample size using other criteria (e.g, assurance probability around NB for the final sample size). Calculations can be done for the entire sample, as well as within subgroups, for example, imposing fairness criteria around the precision of estimates among minority sub-groups.
@@
The rest of this manuscript is structured as follows. First, we briefly review common metrics of model performance and multi-criteria sample size formula by Riley et al. We will then introduce a Bayesian extension of this framework based on various sample size rules. We propose a general approach for characterizing uncertainty around model performance based on commonly reported information from previous studies. We outline Monte Carlo sampling algorithms for drawing from the posterior distributions of relevant quantities. A case study based on a model for predicting COVID-19 deterioration showcases the developments. We conclude by suggesting further areas for research.
2 Methods
@@
We review the context, common metrics of model performance, Riley’s sample size formulas for external validation studies, and our proposal for Bayesian sample size determination.
2.1 Context
@@
We focus on external validation of a model for predicting the risk of a binary outcome. Broadly speaking, an external validation study is an endeavor in learning about the joint distribution of predicted risks ($\pi$) from a pre-specified model and observed outcomes ($Y$: 0 no-event, 1 event) in a target population. The validation sample $D_{N}$ of size $N$ can be seen as $N$ pairs of predicted risks and observed results: $D_{N}=\{(\pi_{i},Y_{i})\}_{i=1}^{N}$.
@@
A classical validation study does not involve learning about the relationship between predictors and the outcome. Rather, the focus is on quantifying the performance of a pre-specified model, and ultimately identifying whether it is worth implementing in the target population. For ease of expositions, we assume one model is being considered. This framework can be extended to multi-model comparisons with relative ease, but we leave this to subsequent exploration. As is the standard in contemporary practice, predicted risks for the model are taken as fixed values, representing the expected outcome probability among all individuals with the same predictor pattern. The proposed approach is applicable for both classical regression models and black-box machine learning algorithms, as a validation study is not concerned on how predictors are used to compute the predicted risk.
2.2 Common metrics of model performance
@@
2) Calibration function $h()$: This is the function that returns the expected actual outcome risk among individuals with a given predicted risk: $h(\pi)=P(Y=1|\pi)$. This function converts the predicted risk for the ith individual, $\pi_{i}$, to the true risk for that person, which we denote by $p_{i}$. $p_{i}:=h(\pi_{i})$. When using non-functional form $h$ we refer to the parameters that define $h()$. Commonly, including in Riley’s sample size formula, $h()$ is modeled linearly on the logit scale: $\mbox{logit}(h(\pi))=\alpha+\beta\mbox{logit}(\pi)$, and thus $h$ consists of an intercept ($\alpha$) and slope ($\beta$). Sometimes studies report $\beta$ and ‘mean calibration’, i.e., $\mathbb{E}(Y-\pi)$, or $\beta$ and observed-to-expected outcome ratio ($O/E:=\mathbb{E}(Y)/\mathbb{E}(\pi)$), but these are equivalent as for a fixed value of $\beta$, knowing any of calibration intercept, mean calibration, or O/E ratio is enough to specify the calibration line. In other instances, $h()$ is modeled more flexibly using non-parametric methods (e.g., based on LOESS smoothing)16. Riley et al, while determining sample size based on a logit-linear calibration function, suggests visually inspecting the variability in anticipated smoothed curves once the sample size is determined2.
@@
4) Net benefit (NB): Details of NB calculations are provided elsewhere17. In a nutshell, at a chosen risk threshold $z$, there are at least three treatment strategies concerning the use of the model: treat no one (with a default NB of 0), use the model to treat individuals whose predicted risk is $\pi\geq z$, or treat all. We index these three strategies by 0, 1, and 2, respectively. The $NB()$ function can be written as (for brevity of notations, in what follows we drop the notation that would indicate NB-related calculations are based on the chosen risk threshold):
@@
2.3 Riley’s sample size formulas
Citing the commonality of the above-mentioned metrics in risk modeling studies, Riley et al structured their multi-criteria equations around desired precision levels (width of 95%CI) around these metrics3. In particular, the following equations were suggested for approximating the standard deviation of the sampling distribution of the estimator (i.e., standard error [SE]), from which Wald-type confidence intervals could be constructed:
@@ 2.3 Riley’s sample size formulas
Assuming all measures are of interest, the final sample size is decided by the largest $N$ among each component. Riley et al proposed also targeting desired CI bands around standardized NB. However, a Bayesian approach facilitates using decision-theoretic approaches for NB, addressing the criticisms around the relevance of conventional inferential methods for NB9,10.
@@
2.4 Going Bayesian
A Bayesian approach towards sample size calculation requires 1) identifying metrics of interest for model performance that will be estimated from this sample (e.g., c-statistic, calibration slope), 2) deciding on precision targets for such metrics (e.g., a 95% CI width of 0.1 for c-statistic), 3) deciding on sample size rules on such metrics (e.g., targeting the expected CI width, demanding a 90% assurance that the CI width will be met, or 90% assurance that the future validation study will correctly identify the most optimal decision at a risk threshold of interest; 4) specifying our current information about model performance; 5) generating random draws from the posterior distribution of desired targets that will be processed according to the specified rules.
@@ 2.4 Going Bayesian
Let $P_{\theta}(\pi,Y)$ represent our knowledge about the joint distribution of predicted risks and outcomes in the target population. This joint distribution is indexed by parameters $\theta$, a random entity summarizing our current knowledge about this joint distribution (which encompasses our beliefs about model performance in the target population). We learn about $\theta$ by collecting ${D}_{N}$. The key aspect of a Bayesian approach is that we treat $\theta$, $D_{N}$, and therefore any quantity (e.g., CI widths) derived from $D_{N}$, as random entities. Implementing this approach involves the following steps: we specify $P(\theta)$, our current (prior) information about model performance in the target population. Repeatedly, we sample from $P(\theta)$ and simulate a random validation sample $P(D_{N}|\theta)$, from which performance metrics are quantified and recorded. Repeating this many times will generate draws from the posterior distribution of these performance metrics for a validation dataset of a particular sample size. These can be summarized in terms of their mean (expected value) and variance, and lead to various sample size rules and assurance probabilities.
Naturally, Bayesian Monte Carlo sampling results in draws from posterior distributions of targets (e.g., CI widths) given a fixed $N$. Sample size calculation thus becomes a stochastic inverse problem: find the smallest $N$ that satisfies a set of sample size rules. As the latter is detached from the core of a Bayesian workflow, we primarily focus on calculating precision and VoI metrics for a given $N$, and discuss our implementation of stochastic root-finding algorithms for sample size determination afterwards.
2.4.1 Sample size rules
@@
2.4.1.1 Sample size rules for metrics of discrimination and calibration
@@
Expected CI widths (eciw): This is related to the Average Length Criterion discussed by Joseph et al13. Here, we target the expected CI width (eciw) across the distribution of ${D}_{N}$:
$eciw(N)=\mathbb{E}_{\mathbf{\theta}}\mathbb{E}_{{D}_{N}|\mathbf{\theta}}(g({D}_{N}))$
If the sample size is fixed, $eciw(N)$ is reported for each component as the expected precision of the future study. For sample size calculation, we seek the minimum $N$ that results in $eciw(N)$, meeting precision target $\tau$: $\min\{N\in\mathbb{N}|eciw(N)\leq\tau\}$. We derive $N$s separately for each component, and choose the maximum $N$ as the final sample size. For example, we might identify the $N$ required to ensure that the expected CI widths of the calibration slope, and c-statistic all meet a desired expected CI width.
Assurance-type rules based on quantiles of CI width (qciw): This is related to the modified worst outcome criterion as discussed by Joseph et al13. Here, we target the probability of meeting (or exceeding) desired precision targets. For a given sample size, this approach returns the CI widths corresponding to a desired quantile $q$ (e.g., $q=0.9$ for 90% assurance):
@@
$F_{N}(q)=P(g(D_{N})\leq q)=\mathbb{E}_{\mathbf{\theta}}\mathbb{E}_{{D}_{N}|\mathbf{\theta}}(I(g({D}_{N})\leq q))$
being the CDF of the distribution of the anticipated CI width for sample size $N$.
Again, for a fixed-$N$ setup we report $qciw(N)$. For sample size determination, we find the minimum $N$ such the the $q^{th}$ quantile is not greater than the target CI width $\tau$: $\min\{N\in\mathbb{N}|qciw(N,q)\leq\tau\}$.
2.4.1.2 Sample size rules for NB
Optimality assurance for NB : This is the expected probability that we will correctly identify the strategy that has the highest population NB based on $D_{N}$. To proceed, let $NB_{D_{N}}(k)$ be the sample estimate of $NB(k)$, and $NB_{\theta}(k)$ its true value given $\theta$. We note that with the future data at hand, the investigator will declare a winning strategy as the one that has the highest expected NB solely based on the sample:
$W(D_{N})=\mbox{arg}\max_{k}(NB_{D_{N}}(k))$
@@ 2.4.1.2 Sample size rules for NB
This assurance is in the probability scale that is non-decreasing and asymptotic to 1 as the sample size is increased.
Expected Value of Sample Information: The key VoI quantity is EVSI(N), the expected gain in NB from conducting a future validation study of size $N$.
@@
The first term on the right hand side is the expcted NB of the decision that we will declare as optimal based on $D_{N}$. The second term on the right-hand side is the expected NB of the best decision under current information. However, it might be the case that without the validation study, the model will not be implemented, regardless of its potential superiority under current information. In which case, this term can be replaced by the NB of the default strategy (e.g., treating no one or treating all).
EVSI(N) is a non-decreasing function. Its maximum value occurs when $N$ is infinity, that is, we have access to the entire population and can unequivocally determine the best strategy. The expected gain from such perfect information is called the expected value of perfect information (EVPI) and can be calculated as (18,19):
@@
Another sample size rule can be based on the expected net benefit of sampling (ENBS)20. ENBS requires scaling the EVSI by the expected size of the population affected by the decision, and subtracting the costs of collecting a sample in the same (true or false positive) unit as NB. Let $\mathcal{M}$ be the expected number of times the decision of interest is to be made, and $\mathcal{W}$ be the effort of every additional recruitment for the validation study in NB units. Then
@@
and the optimal sample size is the one that maximizes ENBS. This rule for sample size determination is fully decision-theoretic as it avoids specifying any arbitrary threshold values. However, it requires scaling total population size and establishing the trade-off between sampling efforts and clinical utility; both of these tasks are context-specific. As such, while we propose this rule here for completeness, we will not investigate it in the case study.
2.4.2 Specifying $P(\theta)$
The Bayesian approach requires specifying $P(\theta)$, the distribution of parameters that govern $P_{\theta}(\pi,Y)$. This parameterization can be done in different ways. For example, if a (pilot) sample form the target population is available, one can adopt a non-parametric approach based on the Bayesian bootstrap. In this scheme, $\theta$ is the vector of weights assigned to each observation in the pilot sample. There weights are random with a distribution of $P(\theta)\sim\mbox{Dirichlet}(1,...,1)$. The empirical distribution of this weighted sample can be considered as a random draw from the distribution of the population21. The validation sample $D_{N}$ then can be obtained from sampling with replacement. Details of this two-level resampling approach is provided elsewhere18,22.
Our focus here is on parametric modeling based on summary statistics from previous studies. Noting that $P(\pi,Y)=P(\pi)P(Y|\pi)$, with $Y|\pi\sim\mbox{Bernoulli}(h(\pi))$, an intuitive way of parameterizing $\theta$ is via specifying $P(\pi)$, the distribution of predicted risks, and $h$, the parameters defining the calibration function $h()$. However, specifying the distribution of predicted risks in the target population is not straightforward, as risk modeling studies seldom provide such information. As well, this specification is not directly aligned with the framework of Riley et al, which demands specifying outcome prevalence, c-statistic, and the calibration function. Fortunately, specification of our knowledge in terms of outcome prevalence, c-statistic, and calibration function, that is, defining $\theta=\{\phi,c,h\}$, can, under mild regularity conditions, fully identify $P(\pi,Y)$. The regularity conditions are as follows:
@@ 2.4.2 Specifying $P(\theta)$
    $P(p)$ is quantile-identifiable; i.e., any two quantiles of the distribution are sufficient for uniquely identifying it. Typical distributions for risks, including Beta, Logit-normal, and Probit-normal satisfy this requirement23.

### v2

@@ Abstract
Contemporary sample size calculations for external validation of risk prediction models require users to specify fixed values of assumed model performance metrics alongside target precision levels (e.g., 95% CI widths). However, due to the finite samples of previous studies, our knowledge of true model performance in the target population is uncertain, and so choosing fixed values represents an incomplete picture. As well, for net benefit (NB) as a measure of clinical utility, the relevance of conventional precision-based inference is doubtful. In this work, we propose a general Bayesian framework for multi-criteria sample size considerations for prediction models for binary outcomes. For statistical metrics of performance (e.g., discrimination and calibration), we propose sample size rules that target desired expected precision or desired assurance probability that the precision criteria will be satisfied. For NB, we propose rules based on Optimality Assurance (the probability that the planned study correctly identifies the optimal strategy) and Value of Information (VoI) analysis. We showcase these developments in a case study on the validation of a risk prediction model for deterioration of hospitalized COVID-19 patients. Compared to the conventional sample size calculation methods, a Bayesian approach requires explicit quantification of uncertainty around model performance, and thereby enables flexible sample size rules based on expected precision, assurance probabilities, and VoI. In our case study, calculations based on VoI for NB suggest considerably lower sample sizes are needed than when focusing on precision of calibration metrics.
footnotetext: From Faculty of Medicine and Faculty of Pharmaceutical Sciences, The University of British Columbia (MS); Department of Statistics, The University of British Columbia (PG); British Columbia Centre For Disease Control (SS); Department of Epidemiology, CAPHRI Care and Public Health Research Institute, Maastricht University, and Department of Development and Regeneration, KU Leuven (LW); Department of Applied Health Sciences, School of Health Sciences, College of Medicine and Health, University of Birmingham, and National Institute for Health and Care Research (NIHR) Birmingham Biomedical Research Centre (RDR)footnotetext: $\dagger$ Co-senior authors with equal contributionfootnotetext: * Correspondence to Mohsen Sadatsafavi, Room 4110, Faculty of Pharmaceutical Sciences, 2405 Wesbrook Mall, Vancouver, BC, V6T1Z3, Canada; email: msafavi@mail.ubc.ca
Introduction
@@
Developing risk prediction models or validating existing ones in a new population represents significant investment in time, resources, and expertise. Like other empirical experiments, the design of such studies should be based on objective, transparent, and defendable principles. A particular aspect of study design is the sample size of such studies. The field has witnessed significant recent developments on this front[1, 2, 3, 4, 5, 6, 7, 8]. An example is the multi-criteria approach by Riley et al for the sample size required for external validation studies, targeting pre-specified widths of the 95% confidence intervals around metrics of model performance[2, 3, 9]. For binary outcomes, such metrics are related to discrimination, calibration, and net benefit (NB). The sample size required for each component is computed separately, with the largest one advertised as the final requirement[3].
@@
In addition to target precision criteria, each component of this approach requires as input an assumed true value of the metric of interest in the target population. In reality, we do not know the true value of model performance metrics with certainty (otherwise there would be no need to conduct a validation study in the first place). Further, due to sampling variability, the precision obtained in one particular dataset may differ substantially from the expected precision for that sample size. Thus, there is more uncertainty to be accounted for than the Riley criteria allow.
@@
A Bayesian approach towards sample size determination allows this uncertainty to be accounted for, which is advantageous on multiple fronts. First, it enables the full use of existing information on model performance at the time the study is designed, rather than forcing the investigator to express their knowledge as the fixed truth. Second, it enables users to choose from different classes of sample size rules, namely those that target the expected values of precision targets (as in Riley criteria), as well as ‘assurance’-type rules for the probability of meeting (or exceeding) precision targets. The latter can be insightful as focusing on expected values alone does not guarantee that the desired precision will be achieved in one particular dataset, and the investigator may want a stronger assurance against this. For example, an investigator that desires an interval width of 0.1 for the c-statistic might not perceive much benefit from the future interval being narrower, but might have a strong preference against not meeting this target. As such, targeting a sample size that will result in, say, a 90% probability that the CI width will be $\leq$ 0.1 might be more appealing than a sample size that targets an expected CI width of 0.1. Finally, when assessing clinical utility, the relevance of precision-based criteria is challenged[10, 11]. A Bayesian approach enables the use of novel, decision-theoretic approaches based on Value of Information (VoI) analysis, which focuses on the expected gain in clinical utility from increasing sample size[12]. Bayesian approaches for power and sample size calculations are mostly investigated in the context of experimental studies that are aimed at interrogating a null hypothesis[13, 14, 15, 16]. However, risk model validation studies are not generally hypothesis-driven. Rather, the focus is on the precise estimation of a variety of metrics of model performance. As such, Bayesian sample size considerations in this context are worth exploring.
@@
Hence, in this article, we propose a Bayesian version of the sample size formula by Riley et al, which enables the investigator to 1) incorporate their uncertainty around assumed model performance in calculations, 2) use sample size rules that target assurance probabilities for meeting chosen criteria, and 3) use rules based on VoI analysis for clinical utility. The proposed framework can be used in two general ways: to quantify the anticipated precision or VoI outcomes from a planned study if the sample size is fixed (as is the case with validation studies that are based on already collected data), or to determine the minimum sample size that achieves pre-specified precision or VoI criteria. This framework can also be used in a hybrid form: determine the sample size based on certain criteria (e.g., targeting CI width for c-statistic and calibration slope), and investigate the consequence of the chosen sample size using other criteria (e.g, assurance probability around NB for the final sample size). Calculations can be done for the entire sample, as well as within subgroups, for example, imposing fairness criteria around the target precision of estimates among minority sub-groups.
@@
The rest of this manuscript is structured as follows. First, we briefly review common metrics of model performance and multi-criteria sample size formula by Riley et al[3]. We will then introduce our Bayesian extension of this framework based on various sample size rules. We propose a general approach for characterizing uncertainty around model performance based on commonly reported information from previous studies, and outline Monte Carlo sampling algorithms for drawing from the corresponding distributions. A case study based on a model for predicting COVID-19 deterioration showcases the developments. We conclude by suggesting further areas for research.
Methods
@@
In this section we review the context, common metrics of model performance, Riley’s sample size formulas for external validation studies, and our proposal for Bayesian sample size determination.
Context
@@
We focus on external validation of a model for predicting the risk of a binary outcome. Broadly speaking, an external validation study is an endeavor in learning about the joint distribution of predicted risks ($\pi$) from a pre-specified model and observed outcomes ($Y$: 0 no-event, 1 event) in a target population. The validation sample $D_{N}$ of size $N$ participants can be seen as $N$ pairs of predicted risks and observed results: $D_{N}=\{(\pi_{i},Y_{i})\}_{i=1}^{N}$.
@@
A validation study does not involve learning about the relationship between predictors and the outcome. Rather, the focus is on quantifying the performance of an existing model. For ease of expositions, we assume one model is being evaluated. This framework can be extended to multi-model comparisons with relative ease, but we leave this to subsequent exploration. As is the standard in contemporary practice, predicted risks for the model are taken as fixed values, representing the expected outcome probability among all individuals with the same predictor pattern. The proposed approach is applicable whenever a model or algorithm provides estimated risks for individuals (regardless of whether it was developed using regression, machine learning or AI-based methods), as a validation study is not concerned on how predictors are used to compute the predicted risk.
Common metrics of model performance
@@
2) Calibration function $h()$: This is the function that returns the expected actual outcome risk among individuals with a given predicted risk: $h(\pi)=P(Y=1|\pi)$. This function converts the predicted risk for the ith individual, $\pi_{i}$, to the true risk for that person, which we denote by $p_{i}$. $p_{i}:=h(\pi_{i})$, which is the expected values of observed outcome values across all individuals with similar predictor patterns. As needed, by $h$ we refer to parameters describing $h()$. Commonly, including in Riley’s sample size formula, $h()$ is modeled linearly on the logit scale: $\mbox{logit}(h(\pi))=\alpha+\beta\mbox{logit}(\pi)$, and thus $h$ consists of an intercept ($\alpha$) and slope ($\beta$). Sometimes studies report $\beta$ and ‘mean calibration’, i.e., $\mathbb{E}(Y-\pi)$, or $\beta$ and observed-to-expected outcome ratio ($O/E:=\mathbb{E}(Y)/\mathbb{E}(\pi)$), but these are equivalent as for a fixed value of $\beta$, knowing any of calibration intercept, mean calibration, or O/E ratio is enough to specify the calibration line. In other instances, $h()$ is modeled more flexibly using non-parametric methods (e.g., based on LOESS smoothing)[17]. Riley et al, while determining sample size based on a logit-linear calibration function, suggests visually inspecting the variability in anticipated smoothed curves once the sample size is determined[2].
@@
4) Net benefit (NB): Details of NB calculations are provided elsewhere[18]. In a nutshell, at a chosen risk threshold $z$, there are at least three treatment strategies concerning the use of the model: treat no one (with a default NB of 0), use the model to treat individuals whose predicted risk is $\pi\geq z$, or treat all. We index these three strategies by 0, 1, and 2, respectively. The $NB()$ function can be written as (for brevity of notations, in what follows we drop the notation that would indicate NB-related calculations are based on the chosen risk threshold):
@@
Riley’s sample size formulas
Citing the commonality of the above-mentioned metrics in risk modeling studies, Riley et al structured their multi-criteria equations around desired precision levels (width of 95%CI) around these metrics[3]. In particular, the following equations were suggested for approximating the standard deviation of the sampling distribution of the estimator (i.e., standard error [SE]), from which Wald-type confidence intervals could be constructed:
@@
Assuming all measures are of interest, the final sample size is decided by the largest $N$ among each component (and by checking whether the subsequent variability of calibration curves is acceptable - as explained above). Riley et al proposed also targeting desired CI bands around standardized NB. However, a Bayesian approach facilitates using decision-theoretic approaches for NB, addressing the criticisms around the relevance of conventional inferential methods for NB[10, 11].
@@
Going Bayesian
A Bayesian approach towards sample size calculation requires 1) identifying metrics of interest for model performance that will be estimated from this sample (e.g., c-statistic, calibration slope); 2) deciding on precision targets for such metrics (e.g., a 95% CI width of 0.1 for c-statistic); 3) deciding on sample size rules on such metrics (e.g., targeting the expected CI width, demanding a 90% assurance that the CI width will not be larger than targeted, or 90% assurance that the future validation study will correctly identify the most optimal decision at a risk threshold of interest); 4) specifying our current information about model performance; 5) generating random draws from the distribution of desired targets that will be processed according to the specified rules.
@@
Let $P_{\theta}(\pi,Y)$ represent our knowledge about the joint distribution of predicted risks and outcomes in the target population. This joint distribution is indexed by parameters $\theta$, a random entity summarizing our current knowledge about this joint distribution (which encompasses our beliefs about model performance in the target population). We plan to learn about $\theta$ by collecting a sample of data $D_{N}$ from $N$ participants in the target population. The key aspect of a Bayesian approach is that we treat $\theta$, $D_{N}$, and therefore any quantity (e.g., CI widths) derived from $D_{N}$, as random entities. Implementing this approach involves the following steps: we specify $P(\theta)$, our current (prior) information about model performance in the target population. Repeatedly, we sample from $P(\theta)$ and simulate a random validation sample $P(D_{N}|\theta)$, from which performance metrics are quantified and recorded. Repeating this many times will generate draws from the prior predictive distributions (henceforth referred to as ‘pre-posterior’ distribution, the anticipated distribution of the quantity of interest after the future sample is procured) of these performance metrics for a validation dataset of a particular sample size. These draws can be processed according to various sample size rules and assurance probabilities.
Naturally, Bayesian Monte Carlo sampling results in draws from pre-posterior distributions of targets (e.g., CI widths) given a fixed $N$. Sample size calculation thus becomes a stochastic inverse problem: find the smallest $N$ that satisfies a set of sample size rules. As the latter is detached from the core of a Bayesian workflow, we primarily focus on calculating precision and VoI metrics for a given $N$, but also briefly discuss our implementation of stochastic root-finding algorithms for sample size determination.
Sample size rules
@@
Sample size rules for metrics of discrimination and calibration
@@
Expected CI widths (ECIW): This is related to the Average Length Criterion discussed by Joseph et al[14]. Here, we target the expected CI width (ECIW) across the distribution of ${D}_{N}$:
$ECIW(N)=\mathbb{E}_{\mathbf{\theta}}\mathbb{E}_{{D}_{N}|\mathbf{\theta}}(g({D}_{N})).$
If the sample size is fixed, $ECIW(N)$ is reported for each component as the expected precision of the future study. For sample size calculation, we seek the minimum $N$ that results in $ECIW(N)$ meeting precision target $\tau$: $\min\{N\in\mathbb{N}|ECIW(N)\leq\tau\}$. We derive $N$s separately for each component, and choose the maximum $N$ as the final sample size. For example, we might identify the $N$ required to ensure that the expected CI widths of the calibration slope and c-statistic all meet a desired expected CI width.
Assurance-type rules based on quantiles of CI width (QCIW): This is related to the modified Worst Outcome Criterion as discussed by Joseph et al[14]. Here, we target the probability of meeting (or exceeding) desired precision targets. For a given sample size, this approach returns the CI widths corresponding to a desired quantile $q$ (e.g., $q=0.9$ for 90% assurance):
@@
$F_{N}(x)=P(g(D_{N})\leq x)=\mathbb{E}_{\mathbf{\theta}}\mathbb{E}_{{D}_{N}|\mathbf{\theta}}(I(g({D}_{N})\leq x))$
being the CDF of the distribution of the anticipated CI width for sample size $N$. Again, for a fixed-$N$ setup we report $QCIW(N)$. For sample size calculation, we find the minimum $N$ such the the $q^{th}$ quantile is not greater than the target CI width $\tau$: $\min\{N\in\mathbb{N}|QCIW(N,q)\leq\tau\}$.
Sample size rules for NB
Optimality Assurance for NB : This is the probability that we will correctly identify the strategy that has the highest population NB based on $D_{N}$. To proceed, let $NB_{D_{N}}(k)$ be the sample estimate of $NB(k)$, and $NB_{\theta}(k)$ its true value given $\theta$. We note that with the future data at hand, the investigator will declare a winning strategy as the one that has the highest expected NB solely based on the sample:
$W(D_{N})=\mbox{arg}\max_{k}(NB_{D_{N}}(k)).$
@@
This assurance is in the probability scale that is non-decreasing and asymptotes to 1 as the sample size is increased.
Expected Value of Sample Information (EVSI): The key VoI quantity is $EVSI(N)$, the expected gain in NB from conducting a future validation study of size $N$, compared with the NB of the decision made with current information:
@@
The first term on the right-hand side is the expcted NB of the decision that we will declare as optimal based on $D_{N}$. The second term on the right-hand side is the expected NB of the best decision under current information. However, it might be the case that without the validation study, the model will not be implemented, regardless of its potential superiority under current information. In which case, this term can be replaced by the NB of the default strategy (e.g., treating no one or treating all).
$EVSI(N)$ is a non-decreasing function. Its maximum value occurs when $N$ is infinity; that is, we have access to the entire population and can unequivocally determine the optimal strategy. The expected gain from such perfect information is called the Expected Value of Perfect Information ($EVPI$) and can be calculated as[19, 20]:
@@
Another sample size rule can be based on the Expected Net Benefit of Sampling ($ENBS$)[21]. $ENBS$ requires scaling the EVSI by the expected size of the population affected by the decision, and subtracting the costs of collecting a sample in the same (true or false positive) unit as NB. Let $\mathcal{M}$ be the expected number of times the decision of interest is to be made, and $\mathcal{W}$ be the effort of every additional recruitment for the validation study in NB units. Then:
@@
and the optimal sample size is the one that maximizes $ENBS$. This rule for sample size determination is fully decision-theoretic as it avoids specifying any arbitrary threshold values. However, it requires scaling NBs to the population and establishing the trade-off between sampling efforts and clinical utility; both of these tasks are context-specific. As such, while we propose this rule here for completeness, we will not investigate it in the case study.
Specifying $P(\theta)$
The Bayesian approach requires specifying $P(\theta)$, the distribution of parameters that govern $P_{\theta}(\pi,Y)$. This parameterization can be done in different ways. For example, if a (pilot) sample form the target population is available, one can adopt a non-parametric approach based on the Bayesian bootstrap. In this scheme, $\theta$ is the vector of weights assigned to each observation in the pilot sample. There weights are random with a distribution of $P(\theta)\sim\mbox{Dirichlet}(1,...,1)$. The empirical distribution of this weighted sample can be considered as a random draw from the distribution of the population[22]. The validation sample $D_{N}$ then can be obtained from sampling with replacement. Details of such two-level resampling approach is provided elsewhere[23, 19].
Our focus here is on parametric modeling based on summary statistics from previous studies. Noting that $P(\pi,Y)=P(\pi)P(Y|\pi)$, with $Y|\pi\sim\mbox{Bernoulli}(h(\pi))$, an intuitive way of parameterizing $\theta$ is via specifying $P(\pi)$, the distribution of predicted risks, and $h$, the parameters defining the calibration function $h()$. However, specifying the distribution of predicted risks in the target population is not straightforward, as risk modeling studies seldom provide such information. Fortunately, specification of our knowledge in terms of outcome prevalence, c-statistic, and calibration function (the same components used in Riley’s equations), that is, defining $\theta=\{\phi,c,h\}$, can, under mild regularity conditions, fully identify $P(\pi,Y)$. The regularity conditions are as follows:
@@
    $P(p)$ is quantile-identifiable; i.e., any two quantiles of the distribution are sufficient for uniquely identifying it. Typical distributions for risks, including Beta, Logit-normal, and Probit-normal satisfy this requirement[24].